OpenStudy (cayla):

Find the slope and standard form of the line that passes through the points (3, 6) and (4, 5).

OpenStudy (cayla):

m=−1,x+y=−9 m=1,x+y=9 m=−1,x+y=7 m=−1,x+y=9

OpenStudy (marypuro):

to find the slope the formula is y^1-y^2/x^1-x^2 so that would come out to 5-6/4-1 which equals -1

OpenStudy (cayla):

@marypuro how to I turn that into standard form

OpenStudy (marypuro):

you would then turn into point slope formula which is y-y^1=m(x-x^1) y=-x+9 and then rewrite x+y=-9